OpenStudy (anonymous):

What percent o 5.92 is 37? would that be 2.8%?

OpenStudy (anonymous):

of *

myininaya (myininaya):

5.92x=37 x=37/5.92=6.25 we want to know what percent 6.25(1)=6.25(100%)=625%

OpenStudy (anonymous):

ohhhh ok. thank u!! :)

myininaya (myininaya):

your welcome just replace the of with multiplication and replace the is with equal and you have a formula for x (or whatever you call the variable)

myininaya (myininaya):

or equation for x